OpenStudy (anonymous):

how does tan(5pi/4) = 1 how does tan(5pi/4) = 1 @Mathematics

OpenStudy (anonymous):

use the unit circle

OpenStudy (anonymous):

a full rotation is 2pi

OpenStudy (anonymous):

i still dont understand sry

OpenStudy (anonymous):

\[\sin\frac{5\pi}{4}=-\frac{\sqrt{2}}{2}\]\[\cos\frac{5\pi}{4}=-\frac{\sqrt{2}}{2}\] \[\tan\frac{5\pi}{4}=\frac{\sin\frac{5\pi}{4}}{\cos\frac{5\pi}{4}}=1\]

OpenStudy (anonymous):

ok thanks joemath

OpenStudy (anonymous):

\[\tan(\frac{5\pi}{4})\] is like \[\frac{\pi}{4} * 5\] so you can just evaluate \[\tan(\frac{\pi}{4})\] and then see what quadrant 5pi/4 is in and then figure out if tan is positive or negative in that quadrant.

OpenStudy (anonymous):

I strongly recommend you memorize the points to \[\frac{\pi}{3}, \frac{\pi}{4}, \frac{\pi}{6}\] because if you know those three then you know any point of any angle.

OpenStudy (anonymous):

alright, i just jot those down, how do i use them to find any point of any angle

OpenStudy (anonymous):

dont forget:\[\frac{\pi}{2}\]No love for the right angles :( lol

OpenStudy (anonymous):

lol yeah.. \[\frac{\pi}{2}, \pi , \frac{3\pi}{2}\] are also good to memorize

OpenStudy (anonymous):

and \[2\pi\]

OpenStudy (anonymous):

but say you want to evaluate \[\sin(\frac{2\pi}{3})\]. This is the same as \[2 * \frac{\pi}{3}\] so just evaluate \[\sin(\frac{\pi}{3})\] which you have memorized. Then figure out what quadrant \[\frac{2\pi}{3}\] is in and then figure out if sin is positive or negative in that quadrant.

OpenStudy (anonymous):

so just by memorizing the points to the angles 30, 45, and 60 you can solve anything

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

hey tyler im glad u came around again...i had two other questions

OpenStudy (anonymous):

first how does the 1st positive two asymptote of y= cot(2x - pi/2) = pi/4 and 3pi/4

OpenStudy (anonymous):

ok.. first you have to know that \[\cot(x) = \frac{1}{\tan(x)}\]

OpenStudy (anonymous):

now in order for a function to have a asymptote, the denominator must equal 0.. So we need to find where tangent is 0

OpenStudy (anonymous):

but the unit circle only has cos and sin...how do i determine tan

OpenStudy (anonymous):

\[\tan(x) = \frac{\sin(x)}{\cos(x)}\]

OpenStudy (anonymous):

so tan would be 0 at 0 and pi

OpenStudy (anonymous):

yeah thats what i was thinking because everything else on the unit circle, if you set it at sin/cos the minimum is like 1

OpenStudy (anonymous):

so using 0 and pi...how do get to pi/4 and 3pi/4

OpenStudy (anonymous):

right.. now you factor out a 2 out of the parenthesis\[y = \cot(x - \frac{\pi}{4})\] to get a phase shift of + pi/4

OpenStudy (anonymous):

so the standard cot(x) functions has a asymptote at 0 and pi but its shifted to the right by pi/4 units.. so add pi/4 to 0 and pi.. but i get pi/4 and 5pi/4.. hmmm

OpenStudy (anonymous):

this is what i have... tanu is undefined when cosu = 0. The first positive number where cosu = 0 is u = pi/2. Solve 2x + pi = pi/2 to find the vertical asymptote. 2x + pi = pi/2 2x = -pi/2 x = -pi/4 -pi/4 is negative of course so try the next number where cosu = 0, 3pi/2. 2x + pi = 3pi/2 2x = pi/2 x = pi/4 You want a second positive asymptote so solve for 5pi/2. 2x + pi = 5pi/2 2x = 3pi/2 x = 3pi/4 The first two positive asymptotes are x = pi/4 and x = 3pi/4.

OpenStudy (anonymous):

hmm you got the right answer.. i dunno where i messed up at :O

OpenStudy (anonymous):

maybe i am going about it wrong.

OpenStudy (anonymous):

ok thats fine...i can email my teacher about that...also how do you determine when sin=1/3 cos= 2/sqrt2 i dont see 1/3 on the unit circle

OpenStudy (anonymous):

There isnt a 1/3 on a unit circle so you have to use a right triangle.

OpenStudy (anonymous):

wait hold on.. lol

OpenStudy (anonymous):

i made a huge mistake lol

OpenStudy (anonymous):